1 1/3 x
1 7/8 x
1 4 /5 =

Answer by Alan3354(69443)

1 1/3 x 1 7/8 x 1 4 /5 =
-------------
= 4/3 * 15/8 * 9/5
= 4*15*9/3*8*5
= 540/120
= 9/2
= 4 1/2